OpenStudy (anonymous):

The volume of a cube is 512 cm3. What would change if the volume of the cube were measured in cubic meters? A. The volume of the prism would decrease. B. The volume of the prism would increase. C. The number of cubic units would increase. D. The number of cubic units would decrease.

OpenStudy (anonymous):

volume remains the same only cubic units will decrease. 1m=100 cm \[1 m^3=(100 cm)^3=(100)^3 cm^3\]